A simple queuing system has the mean interval time of 8 minutes and a mean service time of 4 minutes .

i) Determine the mean service rate and the mean arrival rate.

ii) Determine the traffic intensity.

iii) Determine the mean time a customer spends in the queue and in the system .

iv) What is the expected number of customers in the queue and in the system.

v) What is the probability of having at most four customers in the system.


1
Expert's answer
2021-09-21T11:44:19-0400

Solution:

(i) Mean service rate"=\mu=\dfrac1{4}"

And mean arrival rate"=\lambda=\dfrac1{8}"

(ii) Traffic intensity"=\rho=\dfrac{\lambda}{\mu}=\dfrac{1/8}{1/4}=0.5"

(iii) Number in the queue"=L_q=\dfrac{\rho^2}{1-\rho}=\dfrac{0.5^2}{1-0.5}=0.5"

Wait in the queue"=W_q=L_q/{\lambda}=0.5/(1/8)=4\ mins"

Wait in the system"=W=W_q+1/\mu=4+4=8\ mins"

Number of the system"=L=\lambda W=(1/8)\times 8=1"

Thus, the mean time a customer spends in the queue "=W_q=4\ mins"

And in the system"=W=8\ mins"

(iv) The expected number of customers in the queue "=L_q=0.5"

And in the system "=L=1"

(v)  Probability that there are zero customers or units in the system"=P_0=1-\rho=1-0.5=0.5"

Now, the probability of having at most four customers in the system"=P_0+P_1+P_2+P_3+P_4"

 We know that probability that there are n customers in the system"=P_n=\rho^nP_0"

So, the probability of having at most four customers in the system

"=0.5+0.5(0.5)^1+0.5(0.5)^2+0.5(0.5)^3+0.5(0.5)^4
\\=0.5(1+(0.5)^1+(0.5)^2+(0.5)^3+(0.5)^4)
\\=0.96875"